OpenStudy (anonymous):

simplify completely x^2 - 3x - 54 / x^2 - 18x +81 (multiply) x^2 + 12x + 36 / x + 6

OpenStudy (anonymous):

Its weird because I get the ones with vairables, but these which most of the time would seem easier, just doesnt click and for me seems to take 10 times more time/ effort then the ones with variables.

OpenStudy (anonymous):

\[\Large \frac{x^{2}-3x-54}{x^{2}-18x+81}*\frac{x^{2}+12x+36}{x+6}\] Step 1. Factor the first numerator. Step 2. Factor the first denominator. Step 3. Factor the second numerator. Keep in mind: \[\Large x^{2}+(a+b)x+(a*b)\] In otherwords, what two numbers when added together get you the middle number, and when multiplied get you the last one? Example: Step 1: 6+-9=-3 6*(-9) = -54 Therefore step 1 = (x+6)(x-9)

OpenStudy (ganpat):

(x2 -3x -54) / (x2 -18x + 81) * (x2 +12x +36) / (x+6) x2 -3x -54 = x2 -9x +6x -54 = x(x-9) +6 (x-9) = (x - 9) (x +6) x2 -18x + 81= x2 -9x -9x + 81 = (x - 9) (x -9) x2 +12x +36 = x2 +6x +6x +36 = (x +6) ( x +6) Now substituteall this values in equation..